# The 1979 Iranian Revolution and the Lost Decade of Science: A Counterfactual Scientometric Analysis

### v1

@@ Abstract
This paper presents a comprehensive scientometric analysis of the long-term impact of the 1979 Iranian Revolution on the nation’s scientific development. Using Scopus-indexed data from 1960 to 2024, we benchmark Iran’s publication trajectory against a carefully selected peer group representing diverse development models: established scientific leaders (Netherlands), stable regional powers (Israel), and high-growth ”Asian Tigers” (South Korea, Taiwan, Singapore) alongside Greece and China. The analysis reveals a stark divergence: in the late 1970s, Iran’s scientific output surpassed that of South Korea, China and Taiwan. The revolution, however, precipitated a collapse, followed by a ”lost decade” of stagnation, precisely when its Asian peers began an unprecedented, state-driven ascent. We employ counterfactual models based on pre-revolutionary growth trends to quantify the resulting ”knowledge deficit.” The findings suggest that, in an alternate, stable timeline, Iran’s scientific output could have rivaled South Korea’s today. We further outline a research agenda to analyze normalized impact metrics (such as FWCI) and collaboration patterns, complementing our findings on publication volume. By contextualizing Iran’s unique trajectory, this study contributes to a broader understanding of the divergent recovery patterns exhibited by national scientific systems following profound political shocks, offering insights into the enduring consequences of historical disruptions on the global scientific landscape.
Keywords: Scientometrics; Counterfactual analysis; Iranian Revolution; Knowledge deficit; Scientific growth; Comparative research systems
@@ 1 Introduction
A nation’s scientific output is a robust indicator of its innovation potential, human capital, and long-term economic prospects [9]. While scientific systems often grow organically, their trajectories can be fundamentally reshaped by profound external events. However, the patterns of recovery following such shocks are not uniform and offer crucial insights into the resilience and dynamics of national innovation systems. Historical analysis reveals at least two distinct archetypes of recovery from systemic political shocks.
@@
The first can be described as a model of **prolonged stagnation followed by slow, policy-driven reconstruction**. The collapse of the Soviet Union provides a powerful example; its formidable scientific enterprise endured a fifteen-year period of steady stagnation due to economic collapse and loss of institutional capacity. A recovery phase did not begin until 2006, initiated by new state policies that tied funding to research assessments [7]. The second archetype is one of a **near-total reset followed by explosive, state-led exponential growth**. China’s trajectory after its Cultural Revolution (1966-1976) exemplifies this path. During this period, scientific work effectively came to a halt, with its output in top international journals dropping to nearly zero by 1973. Yet, immediately following the end of the revolution in 1976, its scientific output began a period of sustained, exponential growth, rebuilding from a near-zero baseline [12].
This paper examines a third, defining event—the 1979 Iranian Revolution—and argues that its aftermath reveals a distinct, hybrid pathway. Prior to 1979, Iran was on a path of rapid scientific modernization, establishing universities and research centers as part of a national development strategy [11]. Bibliometric data from this era confirms a nascent system experiencing exponential growth. Most strikingly, Iran had surpassed the scientific output of nations like South Korea and Taiwan, countries that would later become exemplars of knowledge-based economies through targeted, long-term science and technology policies [8].
The revolution, the subsequent Iran-Iraq War (1980–1988), university closures, and international isolation triggered a sharp collapse in scientific productivity [10]. Iran’s subsequent journey—an initial collapse, a long ”lost decade” of stagnation reminiscent of the post-Soviet experience, and then a rapid but incomplete recovery—does not fit neatly into either of the established recovery models. This study situates Iran’s interrupted journey within this comparative theoretical framework to better understand the opportunity cost of the disruption.
@@
Using a scientometric approach, this paper addresses two primary questions: 1) What was the immediate and long-term impact of the 1979 Revolution on Iran’s scientific publication output compared to its peers? 2) What might Iran’s scientific standing be today if its pre-revolutionary growth had continued uninterrupted? We answer the second question by developing counterfactual models to quantify the ”knowledge deficit.” By systematically analyzing the Iranian case, this study moves beyond a national case study to contribute to a broader theory of scientific system resilience. It seeks to explain how different initial conditions—including the nature of the political shock, the level of prior integration into global science, and the subsequent state ideology—lead to divergent pathways of scientific recovery, thereby refining our understanding of the compounding costs of historical disruption.
@@ 2.1 Data Source and Query
The primary data source for this study is the Elsevier Scopus database, which offers comprehensive coverage of peer-reviewed literature. We collected annual publication counts for “articles” and “reviews” for Iran and its seven peer countries (Netherlands, Israel, Greece, China, Singapore, South Korea, Taiwan) for the period 1960–2024.
The data was extracted using Scopus’s advanced search interface. For each country, the query was structured as follows: AFFILCOUNTRY(country_name) AND (DOCTYPE(ar) OR DOCTYPE(re)) AND PUBYEAR > 1959 AND PUBYEAR < 2025
The resulting annual counts form the basis of our time-series analysis. Python 3.12 with Pandas and Matplotlib was used for all analyses.
@@ 2.2 Analytical Approach and Metrics
    • Field-Weighted Citation Impact (FWCI): To measure citation impact relative to the global average in the same field, year, and document type.
    • Outputs in Top 10% Percentiles: To measure the production of elite, highly-cited science.
@@ 2.4 Limitations of the Study
3 Results: A Tale of Two Trajectories - Quantity and Quality
Fig. 1 shows the number of publications of Iran during the period of 1969-2024. Figs. 23 show the same statistics for South Koera and Taiwan.
(a) The foundational era (1960–1979)
(b) The “lost decade” (1980–1999)
(c) The recovery phase (2000–2010)
(d) Plateau and recent trends (2010–2024)
@@ 3 Results: A Tale of Two Trajectories - Quantity and Quality
(a) Foundational era (1960–1979)
(b) Take-off and consolidation (1980–1999)
(c) Scale-up and recent dynamics (2000–2024)
@@
Our analysis reveals a story of divergence that unfolds not only in the volume of scientific output but also in its quality and international impact. We integrate the analysis of publication counts with normalized impact metrics to provide a multi-dimensional view of Iran’s scientific journey.
@@
This period establishes the critical baseline: Iran was not a follower but a leader among its Asian peers. Iran’s output grew exponentially in the 1970s, reaching 518 papers in 1978. In that same year, South Korea produced 260 papers, China has around the same number of papers and Taiwan produced 550. Iran was firmly ahead of South Korea and on par with Taiwan, positioning it as a key emerging scientific player.
@@
This 20-year period is where the paths of Iran and the Asian Tigers irrevocably split. Iran’s output collapsed to a low of 122 papers in 1984 and did not surpass its 1978 peak until 1995. This was its “lost decade.” In stark contrast, this was the breakout period for South Korea and Taiwan. Driven by deliberate, state-led industrial and S&T policy, their publication curves turn sharply upward, embarking on a sustained exponential growth path. By 1999, South Korea was publishing over 17,000 papers and Taiwan over 13,000, while Iran had just reached 1,450. The gap was no longer a gap; it was a chasm. This divergence was not merely quantitative. The international isolation that characterized this period for Iran likely suppressed its citation impact, while the ”Asian Tigers” were rapidly integrating into global scientific networks, a crucial step for improving research quality and visibility.
3.3 2000-2024: Recovery in Scale, Lag in Impact
The final era cemented the new global hierarchy. Iran mounted a remarkable recovery in absolute terms, reaching a peak of $\approx 80,000$ publications in 2022 (Figure 3-c). However, this impressive growth in scale tells only half the story. An analysis of quality metrics reveals a more nuanced picture of an ”unbridgeable gap.”
Figure 4: Annual scholarly output (1960–1995) based on Crossref data. The figure extends historical coverage before Scopus indexing became comprehensive. South Korea and Israel show early exponential scale-up, while Iran’s pre-1979 growth collapses into a “lost decade” during the 1980s, reflecting the stagnation observed in post-revolutionary years.
Figure 5: Approximate year-normalized citation index (yNCI*) for 1960–1995, reconstructed from Crossref’s citation-to-date records. Higher yNCI* values indicate greater relative citation impact. Israel and Greece maintain consistently high values, while Iran’s sharp drop after 1979 confirms the collapse in citation-normalized momentum.
Figure 6: Share of publications in the approximate global top 10% citation percentile (Top-10%*) between 1960–1995. Derived from Crossref data, these results show early volatility and eventual divergence: mature systems (Israel, Netherlands) stabilized above 15–25%, while Iran’s share approached zero during the 1980s, marking the nadir of its research visibility.
@@ 3.4 Pre-Scopus Historical Reconstruction (1960–1995)
To contextualize the Scopus-based results, Figures 46 extend the analysis back to 1960 using Crossref’s metadata and citation-to-date records. Because Scopus coverage before the mid-1990s is incomplete, Crossref provides a valuable approximation of earlier publication and citation trends. These reconstructed metrics—denoted by an asterisk (yNCI*, Top-10%*)—use total citation counts to date rather than fixed five-year windows, offering a reasonable historical baseline for comparing national trajectories prior to the 1979 Revolution. The patterns corroborate the main narrative: Iran’s pre-1979 scientific system exhibited healthy compounding growth and moderate global visibility, both of which collapsed abruptly during the 1980s before partial recovery in the 1990s.
3.5 Trends in Scientific Output and Quality (1996–2024)
Figures 7, 8, and 9 illustrate the evolution of Iran’s scientific performance compared with several peer countries in Asia and Europe based on SciVal data from 1996 to 2024. The three standard field-normalized indicators—Scholarly Output, Field-Weighted Citation Impact (FWCI), and Outputs in Top 10% Percentiles—jointly provide a quantitative picture of both productivity and research quality.
Figure 7: Annual Scholarly Output of Iran and selected countries (1996–2024). The data are based on Elsevier SciVal metrics and include journal articles and reviews only. China exhibits exponential growth, while Iran shows steady post-2000 expansion.
Figure 8: Field-Weighted Citation Impact (FWCI) by publication year (1996–2024). FWCI represents citation impact normalized by field, year, and document type. Iran’s FWCI increased gradually after 2000, approaching the global average (FWCI $\approx$ 1.0) by 2020, while leading countries such as the Netherlands and Singapore consistently remained above 1.5.
Figure 9: Percentage of publications within the Top 10% citation percentiles (1996–2024). The share of Iran’s highly cited papers rose from below 5% in the late 1990s to around 12% by 2024, indicating gradual improvement in research visibility. Advanced systems like the Netherlands and Singapore maintained levels above 25%, showing the gap in global excellence.
The combined trends reveal that while Iran’s research output expanded sharply after the early 2000s, improvements in citation-normalized quality indicators (FWCI and Top 10%) occurred more gradually. The lag between quantity and quality is consistent with the time required for international collaboration and journal visibility to mature.
@@ 3.5 Trends in Scientific Output and Quality (1996–2024)
A similar trend is visible in the production of elite science. Figure 9 shows the percentage of publications in the top 10 percent most-cited papers globally. Iran’s share grew impressively from under 5 percent in the late 1990s to around 12 percent by 2024. Yet, this figure still lags behind top-tier systems like the Netherlands and Singapore, which consistently place over 25 percent of their output in this elite category. This data suggests that while Iran successfully rebuilt the *scale* of its scientific enterprise, it has yet to achieve the same level of global *influence* and *excellence* as the peers it once led.
To synthesize these multi-dimensional findings, we present a comparative profile in Table 1.
@@
3.6 Cross-country synthesis on a log scale (Figure 10)
Figure 10 juxtaposes Iran’s publication trajectory with three peer groupings on a logarithmic scale, so that line slopes approximate growth rates while levels track absolute scale. The vertical dashed line marks the 1979 breakpoint.

### v2

# Iran’s Great Scientific Divergence: Counterfactual Evidence for the Long-Term Shock of the 1979 Revolution

@@ Abstract
This study quantifies the long-term impact of the 1979 Iranian Revolution on Iran’s trajectory of scientific publications, a critical examination for science and public policy. Using comprehensive data from 1960 to 2024, we benchmark Iran against pre-revolutionary peers (e.g., South Korea) and employ multidimensional scientometrics, including the rigorous Synthetic Control Method (SCM). Results demonstrate a significant divergence: Iran, which led its peers in 1978, experienced collapse and stagnation (1980–1999) while peers grew exponentially. SCM, which optimally matches pre-1979 Iran with South Korea, quantifies a cumulative knowledge deficit of approximately 551,000 publications by 2024. Furthermore, despite a post-2000 volume recovery, a persistent quality gap exists; research impact (measured by Field-Weighted Citation Impact) consistently lags behind the global average. This analysis provides a robust, data-driven quantification of the generational opportunity cost of the 1979 disruption on national scientific development.
Keywords: Iranian Revolution; Cultural Revolution; Higher Education; Science Policy; Development; Knowledge Deficit; Counterfactual History; South Korea
@@ 1 Introduction
In the 1960s and 1970s, Pahlavi’s Iran was defined by an ambitious, state-led modernization drive. Central to this national project, often articulated through development plans and concepts like the “Great Civilization” (Tammadon-e Bozorg), was the construction of a modern, Western-aligned scientific and higher education infrastructure. This was a deliberate national development strategy [vakilRevolutionaryEngineers2025]. It fueled the establishment and rapid expansion of elite institutions such as Pahlavi University in Shiraz, designed to be a national beacon, and the Aryamehr University of Technology (now Sharif), intended to train the engineering and technical cadre for a nascent industrial economy. This strategy was predicated on the belief that scientific capacity was synonymous with national sovereignty and long-term economic prosperity [King2004].
The momentum of this pre-revolutionary push was not merely rhetorical. Bibliometric data from this era confirms a nascent system experiencing exponential growth. By the late 1970s, Iran was at a critical takeoff point. In a striking comparison that forms a central theme of this paper, Iran’s total scholarly output had surpassed that of nations like South Korea and Taiwan [Amsden1989]. These countries, which shared a similar trajectory of state-led, Western-aligned development, would later become global exemplars of knowledge-based economies through targeted, long-term science and technology policies. The data from this period establishes a significant baseline: Iran’s pre-1979 scientific trajectory was not only real but was on par with, or even leading, the very peers that would come to define the ”Asian Tiger” model of development.
The 1979 Iranian Revolution, however, fundamentally altered this path. While all revolutions reshape national institutions, the impact on Iran’s scientific enterprise was immediate, profound, and structural. The primary shock was the “Cultural Revolution” (Enqelab-e Farhangi), launched in 1980. With the stated goal of ”Islamizing” universities and purging secular and Western influences, the policy resulted in the complete closure of all higher education institutions from 1980 to 1983. This three-year shutdown was not a pause; it was a systemic reset that severely reduced the nation’s human capital. This event was the first and most direct cause of the collapse in Iran’s scientific growth. As a direct result of the purges, ideological vetting, and forced exoduses that accompanied this period, the number of university teaching staff was nearly halved. According to data from the Ministry of Culture and Higher Education, the total number of teaching staff (including full-time, part-time, and sessional) fell from 16,877 in the 1979-80 academic year to 9,042 by 1982-83, creating a shortage of over 7,800 faculty members [Mojab1991].
This internal, structural shock to the academic system was immediately exacerbated by a profound external crisis: the outbreak of the Iran-Iraq War (1980–1988). The war diverted all national resources and focus toward military defense and national survival. This paper aims to provide the first quantitative analysis of this combination of shocks—university closures, the loss of half the nation’s professoriate, total war, and international isolation—which historical analysis suggests triggered a sharp decline in scientific productivity. We will empirically investigate the period from 1980 through the late 1990s, *termed here the ”lost decade”*, to *test the hypothesis* that it constituted a period of deep stagnation and disconnection from the global scientific community.
@@
Iran’s subsequent journey—an initial collapse, a long stagnation, and then a rapid but incomplete recovery from the late 1990s onward[Stone2005],[hamdhaidari2008higher][mozafari2016iran][golkar2017politics]—does not fit neatly into the established models of scientific system recovery. Historical analysis reveals at least two distinct archetypes of recovery from systemic political shocks. The first can be described as a model of prolonged stagnation followed by slow, policy-driven reconstruction. The collapse of the Soviet Union provides a significant example; its formidable scientific enterprise endured a fifteen-year period of steady stagnation due to economic collapse and loss of institutional capacity. A recovery phase did not begin until 2006, initiated by new state policies that tied funding to research assessments [Yegorov2009]. The second archetype is one of a near-total reset followed by rapid, state-led exponential growth. China’s trajectory after its Cultural Revolution (1966-1976) exemplifies this path. During this period, scientific work effectively came to a halt, with its output in top international journals dropping to nearly zero by 1973. Yet, immediately following the end of the revolution in 1976, its scientific output began a period of sustained, exponential growth, rebuilding from a near-zero baseline [Deng1997]. Iran’s story, with its unique combination of a post-Soviet-style stagnation followed by a rapid, quantitative-focused resurgence, represents a distinct, hybrid pathway.
While historical narratives have captured the political dimensions of the revolution and the war, the long-term opportunity cost of this disruption on Iran’s developmental trajectory remains largely unquantified. This paper situates Iran’s interrupted journey within this comparative theoretical framework to better understand that cost. Using a scientometric approach, this paper moves beyond anecdote to provide a data-driven analysis of this historical divergence. It addresses two primary questions: 1) What was the immediate and long-term impact of the 1979 Revolution on Iran’s scientific publication output compared to its peers? 2) What might Iran’s scientific standing be today if its pre-revolutionary growth had continued uninterrupted?
@@
We answer the second question by developing counterfactual models to quantify the ”knowledge deficit.” While a nation’s scientific output is a robust indicator of its innovation potential and human capital [King2004], this study is not merely a statistical exercise. By systematically analyzing the Iranian case against its peers, this paper provides a new, quantitative lens on the long-term developmental consequences of the 1979 Revolution. It seeks to explain how different initial conditions—including the nature of the political shock, the level of prior integration into global science, and the subsequent state ideology—lead to divergent pathways of scientific recovery, thereby refining our understanding of the compounding costs of historical disruption.
@@ 2.1 Data Source and Query
The primary data source for this study is the Elsevier Scopus database, a comprehensive index of scholarly publications. We collected annual publication counts (limited to documents classified as “articles” or “reviews”) for the period 1960–2024 for Iran and a carefully selected peer group.
This comparative group was not chosen arbitrarily. It was designed to benchmark Iran’s unique trajectory against diverse and relevant developmental models. The peer group includes:
  • High-Growth “Asian Tigers” (South Korea, Taiwan, Singapore): These nations serve as the most critical counterfactuals. Like pre-revolutionary Iran, they were state-led, Western-aligned “developmental states” that successfully executed long-term, strategic visions for building knowledge-based economies.
  • Stable Regional & Mature Powers (Israel, Netherlands): Israel provides a benchmark for a stable, high-performing regional scientific power, while the Netherlands represents a mature, established global scientific leader.
  • Other Comparative Cases (Greece, China): Greece offers a European comparison with a different economic and integration model. China represents the archetype of a near-total systemic reset (after its Cultural Revolution) followed by explosive, state-led growth, providing a different recovery model.
The resulting annual publication counts form the basis of our longitudinal time-series analysis.
@@ 2.2 Analytical Approach and Metrics
    • Field-Weighted Citation Impact (FWCI): To assess research quality and influence, we used the Field-Weighted Citation Impact (FWCI). This standard metric measures citation impact relative to the global average in the same field.
    • Outputs in Top 10% Percentiles: To measure the production of elite, highly-cited science, we analyzed the ’Outputs in Top 10% Percentiles,’ which represents the share of a country’s papers that are among the top 10% most cited globally.
@@ 2.4 Limitations of the Study
3 Results: A Tale of Two Trajectories – Quantity and Quality
Fig. 1 shows the number of publications of Iran during the period 1969–2024. Figs. 2–4 show the same statistics for South Korea, Taiwan, and Turkey.
@@ 3 Results: A Tale of Two Trajectories – Quantity and Quality
Figure 4: Annual Scopus-indexed articles and reviews for Turkey (1960–1979, 1980–1999, and 2000–2024).
Our analysis reveals a story of divergence that unfolds not only in the volume of scientific output but also in its quality and international impact. We integrate publication counts with normalized citation and impact indicators to provide a multidimensional view of Iran’s scientific path in comparison with its Asian peers.
@@
This period establishes the significant baseline: Iran was not a follower but an early leader among its Asian peers. Iran’s output grew rapidly in the 1970s, reaching 518 papers in 1978. In the same year:
- South Korea produced 260 papers, - China was at a comparable level, roughly in the 200–300 range, - Taiwan produced about 550 papers, - Turkey published approximately 470 papers by 1979.
Thus, Iran was clearly ahead of South Korea and roughly on par with Taiwan and Turkey, positioning it as a rising scientific player with strong upward momentum before 1979.
@@
This 20-year period is where the trajectories of Iran and the Asian Tigers permanently diverged.
Iran’s output collapsed to a low of 122 papers in 1984 and did not surpass its 1978 peak until 1995. This period represents Iran’s “lost decade,” reflecting the combined effects of the Cultural Revolution’s university shutdowns and the reallocation of resources following the Iran–Iraq War.
In sharp contrast, South Korea, Taiwan, and China entered their take-off phase:
- South Korea grew from a few hundred papers in the early 1980s to over 17,000 papers in 1999, - Taiwan expanded to more than 13,000 papers, - China began a strong upward trajectory in the 1990s, surpassing 20,000 publications by the late decade, - Turkey also accelerated, rising from a few hundred papers in the early 1980s to nearly 7,000 papers in 1999.
By the end of the century, the gap between Iran and its peers had evolved from a moderate separation into a wide, structural gap. This divergence was not only quantitative: the Asian Tigers were rapidly integrating into global scientific networks—boosting visibility, citation impact, and international collaboration—while Iran was facing geopolitical isolation and limited access to these networks.
3.3 2000–2024: Recovery in Scale, Lag in Impact
In the 21st century, Iran mounted a remarkable recovery in absolute output. Iran grew from about 2,500 papers in 2000 to a peak of roughly 80,000 papers in 2022. This rapid expansion reflects major national investments in mass higher education and research infrastructure.
However, Iran’s surge in quantity does not fully translate into gains in scientific impact.
Meanwhile, its peers followed different trajectories:
- China experienced an unprecedented rise, reaching more than 1,000,000 annual publications by 2024. - South Korea surpassed 100,000 papers annually in recent years. - Taiwan stabilized in the range of 50,000–60,000 papers. - Turkey grew rapidly from about 7,000 papers in 2000 to more than 60,000 papers by 2024.
Despite Iran’s strong quantitative return, citation-based indicators show a persistent lag relative to the Asian Tigers. While South Korea, Taiwan, and Turkey have improved their field-normalized citation impact over time—and China has achieved near-parity with global averages in many fields—Iran remains significantly below these benchmarks. This pattern suggests a decoupling between scale and impact: Iran has achieved mass production, but its influence within the global scientific system expands at a much slower rate.
@@ 3.4 Pre-Scopus Historical Reconstruction (1960–1995)
To contextualize the Scopus-based results, Figures 5, 6 and 7 extend the analysis back to 1960 using Crossref’s metadata and citation-to-date records. Because Scopus coverage before the mid-1990s is incomplete, Crossref provides a valuable approximation of earlier publication and citation trends.
Figure 5: Annual scholarly output (1960–1995) based on Crossref data.
Figure 6: Approximate year-normalized citation index (yNCI*) for 1960–1995, reconstructed from Crossref citation records.
Figure 7: Share of publications in the approximate global top 10%* citation percentile (Top-10%*) between 1960–1995.
These reconstructed metrics—denoted by an asterisk (yNCI*, Top-10%*)-use total citation counts to date rather than fixed five-year windows, offering a reasonable historical baseline for comparing national trajectories prior to the 1979 Revolution. The patterns corroborate the main narrative: Iran’s pre-1979 scientific system exhibited healthy compounding growth and moderate global visibility, both of which collapsed abruptly during the 1980s before partial recovery in the 1990s.
3.5 The Paradox of Quantity vs. Quality
Figures 8, 9 and 10 illustrate the central paradox of Iran’s post-2000 recovery: a successful restoration of research scale (quantity) alongside a persistent lag in research influence (quality).
Figure 8: Annual scholarly output of Iran and selected countries (1996–2024).
Figure 9: Field-Weighted Citation Impact (FWCI) by publication year (1996–2024).
Figure 10: Percentage of publications within the Top 10% citation percentiles (1996–2024).
The combined trends in Figure 8 reveal that Iran’s research output expanded sharply after the early 2000s, demonstrating a remarkable recovery in absolute numbers. However, this growth in scale obscures a significant quality deficit when analyzed with normalized metrics.
@@
A similar trend is visible in the production of elite science. Figure 10 shows the percentage of publications in the top 10 percent most-cited papers globally. Iran’s share grew impressively from under 5 percent in the late 1990s to around 12 percent by 2024. Yet, this figure still lags behind top-tier systems like the Netherlands and Singapore, which consistently place over 25 percent of their output in this elite category.
This persistent gap between quantity and quality, while common in rapidly expanding systems, is particularly pronounced in the Iranian case. We argue this lag is not a temporary phase but the result of enduring structural factors, operating at both the international and domestic levels.
Internationally, the mechanisms are clear. Decades of relative ”isolation, exacerbated by geopolitical constraints and sanctions”, have created a structural limitations on Iranian science. This extends beyond just constraining travel or funding; it has limited opportunities for ”deep collaboration with leading global research groups”. As international collaboration is a primary driver of citation impact (FWCI), this ”reduced integration into elite networks inherently limits knowledge transfer and citation visibility”. This isolation has meant limited access to the latest equipment, shared data, and the informal knowledge transfer that drives scientific frontiers, thereby constraining the potential impact of the research being produced.
Domestically, the drivers of the quality gap are equally significant and are, paradoxically, linked to the quantitative recovery itself. The state-led push to rebuild the scientific enterprise from the late 1990s onward was heavily influenced by a desire to reclaim national prestige through global university rankings. This created a ”science policy and university promotion criteria” system that ”often prioritized quantitative metrics over impact”. This incentive structure created a rational response: ”volume maximization strategies”. This includes ”publishing in lower-tier venues or engaging in incremental research” to meet promotion quotas, often ”at the expense of the long-term, resource-intensive projects that typically yield high-impact breakthroughs”.
Consequently, the very policies that fueled the impressive recovery in scale (Figure 8) may have inadvertently entrenched the deficit in influence (Figures 9 and 10). To synthesize these multi-dimensional findings, we present a comparative profile in Table 1.
@@
3.6 Cross-country synthesis on a log scale
The comparative analysis in Figure 11, which juxtaposes trajectories on a logarithmic scale, provides the clearest visualization of the divergence. It reveals how small differences in growth rates (slopes), when sustained over decades, compound into massive, permanent differences in level. The vertical dashed line marks the 1979 breakpoint.